OpenStudy (anonymous):

two angles are complementary. The sum of the measure of the first angle and half the second angle is 69.5 Find the measure of the angles.

OpenStudy (anonymous):

x+y=90 so x=90-y x+.5y=69.5 plug in 90-y for x in x+.5y=69.5 90-y+.5y=69.5 subtract 90 from both sides and combine the ys -.5y=-20.5 divide both sides by -.5 to isolate the y *** y=41 then plug that value into the first equation x+41=90 subtract 41 from both sides *** x=49

OpenStudy (anonymous):

Let x=first angle then 90-x is the second angle set up equation using given: x+.5(90-x)=69.5 x+45+.5x=69.5 1.5x=24.5 x=49 degrees so the complementary angle must be 90-49= 41 degrees